OpenStudy (anonymous):

Help Please!! In the right triangle ABC, angle A and angle B are complementary angles. The value of cosA is 5/13. What is the value of sin B.

OpenStudy (wolf1728):

 cos A = adj/hyp = 5 / 13 sin B = opp/hyp = 5/13

OpenStudy (anonymous):

thanks a lot